OpenStudy (anonymous):

simplify. 8 sqrt 3 (2 sqrt 6 - 5 sqrt 15)

OpenStudy (angela210793):

is it \[\sqrt[8]{3*(2\sqrt{6}}-5\sqrt{15)}\]

OpenStudy (anonymous):

yeah but the first sqrt sign shouldnt be so big covering all of the problem. it should only cover the 3

OpenStudy (angela210793):

oh i got it :)

OpenStudy (anonymous):

ans is 48root2 - 40root15

OpenStudy (anonymous):

harkirat is right , again :)

OpenStudy (angela210793):

yea

OpenStudy (anonymous):

its ighter choices A) 30-39 sqrt 5 B)16 sqrt 3 -40 sqrt 15

OpenStudy (anonymous):

C)48 sqrt 2 - 120 sqrt 5

OpenStudy (anonymous):

D) 90-3 sqrt 15

OpenStudy (anonymous):

those are the choices they give me

OpenStudy (anonymous):

48root2 - 120root5 is the correct ans

OpenStudy (amistre64):

\[8 \sqrt{3} * (2 \sqrt{6} - 5\sqrt{15})\] \[16\sqrt{18}-40\sqrt{45}\] 18 = 9.2 45 = 9.5 \[54\sqrt{2}-120\sqrt{5}\]

OpenStudy (amistre64):

yeah; 16.3 = 48 .... not 54 lol